Fermi National Laboratory
 

Operation Summary for April 2, 2010 - April 5, 2010

 

The Integrated luminosity for the period from 3/29/10 to 4/5/10 was 71.71 pb-1.
NuMI reported receiving 8.68E18 protons on target during this same period.

Friday April 2

The day shift began with Operations monitoring the Tevatron (TeV) store 7722 (69.86E30), the Recycler stash (325.9E10), and the Antiproton Source stacking (19.9E10), with the Linac, Booster, and Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations reported I- Source regulation problems early on shift. Operations terminated store 7722 at 12:15 PM. Cryo system technicians began maintenance work at 1:48 PM. Operations put the TeV into shot setup and established store 7724 at 2:59 PM with an average initial luminosity (AIL) of 364.7E30. In the early afternoon, a MI RF station (MIRF1) started tripping repeatedly and eventually required a modulator change out.

During the evening shift, Operations reported a couple more MIRF station trips during the shift.

Saturday April 3

The midnight shift began with Operations monitoring store 7724 (114.27E30) and the stash (234.7E10), with Pbar stacking (12.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported a high LCW differential pressure at TeV sector A2 house; Water group personnel came in and fixed the problem.

During the day shift, Operations terminated store 7724 and put the TeV into shot setup at 8:16 AM. Operations established store 7725 at 9:27 AM with an AIL of 342.5E30. Operations established beam to the Meson MTest (T1004) experiment at 10:11 AM.

During the evening shift, Operations disabled the MTest beam at 7:10 PM. The NuMI beam permit tripped off when the output of a power supply (E:V108) drifted low.

Sunday April 4

The midnight shift began with Operations monitoring store 7725 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 7725 at 1:54 AM and put the TeV into shot setup at 2:05 AM. Operations established store 7726 at 2:58 AM with an AIL of 343.8E30. Operations reported that MiniBooNE was off for about fifteen minutes due to an interlocked air bit that spiked too high, pulling its beam permit; it reset Also, three Linac quadrupole power supplies tripped off at the same time and caused a short downtime; they all reset.

During the day shift, Operations established beam to MTest at 8:39 AM.

During the evening shift, Operations disabled beam to MTest at 6 PM. Operations terminated store 7726 at 7:20 PM and put the TeV into shot setup and established store 7727 at 8:39 PM with an AIL of 323.7E30.

Monday April 6

The midnight shift began with Operations monitoring store 7727 and the stash, with Pbar stacking, with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported a short downtime caused by several vacuum pumps that tripped off in the P1 transfer line. The MiniBooNE horn tripped off; it reset. At 3:54 AM, a Linac quadrupole power supply tripped off on a temperature indication. Operators called in Linac techs and they had the power supply repaired and beam back by 6:02 AM. Operators tuned the Booster extraction to improve beam positions.

The Plans for Monday

The plans for today are stack, stash, and store. The Run Coordinator will allow the following work: Linac personnel will flush the I- Source water resistor; TeV personnel will tune their chromaticity tracker; MI personnel will conduct a study; and Recycler personnel will continue to make improvements to their RF system.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/05/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory